1) Explain the following: resolution, reference voltage, quantization error, and step size 2) What is the resolution of the Arduino Uno’s analog to digital converter? Provide the relationship between resolution and step size 3) Convert the following voltages into their corresponding values: 1.33V; 2.71V; 2.75V; 3.99V 4) Convert the following values into their corresponding voltages: 125; 400; 938; 20 5) Plot (line graph) the relationship between the angle of a potentiometer (as described in class and textbook) on the x‐axis and the raw reading of the ADC on the y‐axis. Derive the equation. 6) Connect a thermistor as shown in part 1 below. By following the instructions through part c, write a program to read the ADC channel connected to the thermistor and print the RAW ADC value to the terminal program.(To be done at home as part of the prelab. Demo at lab start)
Answer #1

solution:

1

Resolution:

Resolution can be communicated in two different ways:

i. It is the proportion between the greatest sign estimated to the littlest part that can be settled - typically with a simple to-advanced (A/D) converter.

ii. It is how much a change can be hypothetically recognized, normally communicated as various bits. This relates the quantity of bits of goals to the genuine voltage estimations.

Reference Voltage:

The reference voltage (for the most part known as Vref) is generally the greatest voltage esteem that the D/A converter can reach. This worth relies upon what is associated with the Vref stick.

Quantization error:

Quantization error is the distinction between the simple sign and the nearest accessible computerized an incentive at each examining moment from the A/D converter. Quantization error additionally presents commotion, called quantization clamor, to the example signal. The higher the goals of the A/D converter, the lower the quantization error and the littler the quantization clamor.

Step size:

Step size is the base change in info voltage which can be settled by the ADC. The idea of step size is intently connected with the goals of ADC

2.

The Arduino Uno ADC is of 10 piece goals (so the number qualities from (0-(2^10) 1023)). This implies it will guide input voltages somewhere in the range of 0 and 5 volts into whole number qualities somewhere in the range of 0 and 1023. So for each (5/1024= 4.9mV) per unit.

3. conversion of voltage into value

i. value=(1023/5V)*Voltage

value= (1023/5V)*1.33V

value=272

ii. value= (1023/5V)*2.71V

value=554

iii. value=(1023/5V)*2.75V

value=563

iv. value=(1023/5V)*3.99V

value=816

4. conversion of values into voltages

voltage= (value*5)/1023

i. voltage= (125*5)/1023

voltage=0.61V

ii. voltage= (400*5)/1023

voltage=1.95V

iii. voltage= (938*5)/1023

voltage=4.58V

iv. voltage= (20*5)/1023

voltage=0.097V